# How does online shopping affect offline price sensitivity?

### v1

@@ 3.1 Setting
We study consumers’ offline purchase behavior and price sensitivity using data from a leading pet supplies retailer in Brazil. This firm operates in a multichannel environment, with over 200 brick-and-mortar stores nationwide and a digital presence through both a mobile app and an e-commerce website. The retailer offers a broad selection of pet-related products, including food, medicines, hygiene and grooming items, toys, and accessories. To track consumer purchases across online and offline channels, the retailer uses Brazil’s National Identification Number, which provides a unique identifier for each customer. Online purchases - whether through the app or website - require customers to log into their accounts, making all such transactions traceable to individual IDs. For brick-and-mortar purchases, customers provide their ID at checkout, although it is not mandatory. According to the retailer, around a quarter of offline transactions are not associated with a customer ID, and this rate has remained stable throughout the study period.
@@ 3.2 Dataset Construction
Selection of Brands and Stock Keeping Units (SKUs): Each category contains a large number of pack sizes and SKUs per brand. First, within each category, we focus on specific pack size(s) to ensure that the price variation for each brand is captured accurately – this avoids the need to aggregate price movements across a number of different pack sizes. Focusing on a narrow set of pack sizes to study brand choice within a category is common in the brand choice literature (e.g. Allenby and Rossi (1998); Kim et al. (2002); Chintagunta et al. (2005); Dubé et al. (2010)). The pack sizes we select for each category are as follows: 15KG for dry dog food, 10 KG for dry cat food, 80cm $\times$ 60cm 30-piece mats for dog hygiene, and silica cat sand (1.6KG, 1.8KG, and 2.0KG) for cat hygiene. We provide a detailed discussion of the reasons for selecting these pack sizes in Web Appendix §$\S$§A.1.1.
Next, we aggregate SKUs to the brand level and estimate consumer preferences and price elasticities at the brand level for each product category, similar to Hitsch et al. (2021); Gordon et al. (2013). Specifically, we follow a two-step process for each category. First, for each category, we select the top-selling brands and SKUs that are comparable in usage and are substitutes for each other. Next, we aggregate prices and costs (our instrumental variable to deal with price endogeneity) to the brand level from the SKU level, following the approach described in Hitsch et al. (2021). We provide details on both these steps in Web Appendix §$\S$§A.1.
Selection of Customers: For each category, we restrict our sample to only those consumers who only purchase the specific pack size of interest for that category (as defined earlier), without switching to any other pack sizes within the same category. Next, given that the goal of our study is to investigate the effect of consumers’ adoption of online shopping111“Adoption of online shopping” refers to a customer making his/her first online purchase with the focal retailer in any category, and not just one of the four focal categories. on their offline price sensitivity, we need to ensure that we have sufficiently long pre- and post-adoption periods for each adopter in our sample to effectively measure price elasticities. Therefore, we focus on customers who adopt online shopping in the middle of our observation period, i.e., July 1st, 2019 – June 30th, 2020. So our 30-month data period is divided into three phases: a pre-adoption period (January 1st, 2019 – June 30th, 2019) when all adopters are in their pre-adoption phase, an adoption period (July 1st, 2019 – June 30th, 2020) during which time the adopters in our sample make their first online purchase, and a post-adoption period (July 1st, 2020 - June 30th, 2021), when all the adopters in our sample are in their post-adoption phase. To clarify, a given adopter’s pre- and post-adoption phases are defined as the months before and after they make their first online purchase. Non-adopters only shop offline through the entire data period.
@@
Estimation Sample Construction: For each customer in a given category (both adopters and non-adopters), we organize the data into a panel that captures purchase occasions and brand choices for individual customers over time. Specifically, we define purchase occasions at the customer-month level. If a customer visits a physical store of the retailer during a month and purchases in any category (not limited to the categories we focus on in this study), that customer-month is included as a purchase occasion for all four categories. If a customer makes no purchases in any category during the month, that customer-month is not included as a purchase occasion. In other words, a purchase occasion represents any customer-month during which we observe any offline interaction with the retailer (similar to week-store visit in Gordon et al. (2013))  [^2] Based on our definition of purchase occasion, if a customer visits multiple times in a month, it is treated as one purchase occasion. In our data, customers, by and large, do not purchase from more than one store in a month. More than 96% of customer-month-level purchase occasions involve purchases from a single store. When we process the choice data for each category, for the minor fraction of purchase occasions that involve more than one store, we label these customer-month purchase occasions with the store where the customer spent the most within that category during the month so that each customer-month level purchase occasion is associated with a single store.. For each purchase occasion, we identify the brand chosen by the consumer as the brand which has positive spend.  [^3] We rarely observe customer-month purchase occasions that involve a purchase of more than one brand. All product categories have fewer than 0.86% of total purchase occasions (e.g., including those with multiple visits to a single store in a month, visits to multiple stores and a single visit) with multiple brand purchases. For these few purchase occasions with multiple brand purchases, we assign the brand with the largest spend for that customer during the month, similar to previous brand choice literature, e.g. Erdem and Keane (1996). Some researchers have handled this issue differently, for instance, by having each brand chosen count as an independent single-brand purchase occasion (Krishnamurthi and Raj 1988; Chib et al. 2004). Since the fraction of purchases is so small, we do not think this decision makes a material impact on our results. If the customer does not have positive spend on any brand during a purchase occasion, we label it as choosing the outside option (i.e., “no-purchase”).
@@ 3.3 Summary Statistics
First, we find that there are systematic differences in the demographics (age, income, and gender) of adopters and non-adopters across all four categories. Across all categories, women tend to adopt online shopping more than men. In addition, online adopters tend to be younger and higher-income households. This makes intuitive sense because younger and higher-income consumers are often more familiar with online technology and seek time-saving and convenient shopping options. Details of these demographic characteristics are shown in Web Appendix A.2.
@@ 4 Descriptive Analysis
We now present some descriptive analysis that examines if/how the adoption of online shopping changes consumer behavior during their offline shopping trips in the post-adoption period. Specifically, we focus on two variables that can capture changes in within-customer offline price sensitivity:
@@ 4.1 Price Paid
$\mathcal{I}\{\textrm{Adopter}_{i}\}\times\mathcal{I}\{t\geq T_{i}^{\textrm{adopt}}\}$0.0102 -0.0106∗∗ -0.0046 -0.0160
@@
We now examine how the prices paid by adopters change after online adoption (compared to non-adopters). To that end, we focus on all offline purchase orders (made by adopters and non-adopters) that contain a focal category purchase and estimate the following two-way fixed-effects regression for each category at the customer-order level:
[…]
@@ 4.2 Category Purchase Frequency
$\mathcal{I}\{\textrm{Adopter}_{i}\}\times\mathcal{I}\{t\geq T_{i}^{\textrm{adopt}}\}$0.0751∗∗∗ -0.0497∗∗∗ -0.0617∗∗∗ -0.0654∗∗∗
@@ 5 Firm’s Problem and Overview of Empirical Strategy
The primary objective of the firm is to quantify the causal effect of consumers’ adoption of online shopping on their offline price sensitivity, compared to consumers who shop exclusively offline. However, quantifying this impact presents both methodological challenges and limited managerial relevance for several reasons.
First, adopting the online channel is a consumer-level decision, and firms cannot randomly assign consumers to adopt the online channel or remain offline-only through field experiments. This makes it impossible to directly measure the Average Treatment Effect (ATE) of online adoption on consumers’ offline price sensitivity. Second, although firms could implement an intent-to-treat (ITT) design - such as randomly offering incentives to some offline-only consumers to encourage their online adoption - this approach still does not give us ATE given that some consumers would never adopt online  [^5] We do not obtain ATE under the ITT design, since the causal quantity obtained from 2SLS estimators represents the Average Causal Response, which is a weighted sum of individual treatment effects where weights are compliance scores; For more details, see Angrist and Imbens (1995); Syrgkanis et al. (2019); Mummalaneni et al. (2023).. Third, even if a firm could somehow identify a significant effect of adopting online shopping on consumers’ offline price sensitivity, it cannot force consumers to choose a specific channel. Therefore, obtaining the ATE is not only impractical but also lacks actionable implications. Instead, the main quantity of interest for the firm is the change in offline price sensitivity among consumers who adopt online shopping compared to those who continue to shop exclusively offline i.e. the Average Treatment Effect on the Treated (ATT). Measuring this would help firms adjust their offline pricing strategies in response to their customers adopting online shopping. Thus, our goal is to identify the ATT of online adoption on offline price sensitivity.
@@
    First, we measure category-level own-price elasticities at the customer-month level: In §$\S$6, we use a logit brand choice model where each consumer’s price sensitivity coefficient varies with the extent of their online spending and include customer-brand fixed effects. To address price endogeneity, we use the control functions approach (Petrin and Train 2010), using wholesale cost as an instrument. To efficiently estimate the model given a large number of customer-brand fixed effects, we implement a Minorization-Maximization (MM) algorithm for the logit model (Chen et al. 2021).
@@
    Next, we use the estimated elasticities in the first step to measure the change in offline price elasticities due to online adoption. In §$\S$7, we use a staggered difference-in-differences approach (Callaway and Sant’Anna 2021) to estimate the ATT of online adoption on offline price elasticities after accounting for differences in the timing of adoption between different adopter cohorts.
@@ 6.1 Choice Model and Utility Specification
For each category, we use a logit model of brand choice (Guadagni and Little 1983; Chintagunta et al. 1991) to estimate offline price elasticities at customer-brand-month-level. Suppose that for a given category, there are $J$ brands indexed by $j=1,2,\ldots,J$ plus an outside option indexed by 0. The utility that consumer $i$ obtains from choosing brand $j$ during offline purchase occasion $t$ is given by  [^6] Here, we use a single index $t$ to denote a purchase occasion, which contains both the time (YearMonth) and store information associated with the purchase occasion.:
@@
$+\alpha_{3}\mathcal{I}\{\textrm{BuyLastMonth}_{ijt}\}+\delta_{ij}+\epsilon_{ijt}~{}~{}\forall j\in\{1,\ldots,J\}$
@@
6.2 Price Endogeneity
A common challenge in estimating price elasticities with observational choice data is price endogeneity. This issue arises when the error term, $\epsilon_{ijt}$, contains some time-varying unobserved demand shocks that may be correlated with prices; e.g., if promotions or advertising are correlated with prices. To address this, we apply the control function approach for two endogenous variables, $\textrm{Price}_{jt}$ and $\textrm{Price}_{jt}\times\textrm{cumPercOnlineSpend}_{it}$ (Petrin and Train 2010; Wooldridge 2015). We use the weighted wholesale cost indices for each brand $j$ (constructed as described in §$\S$3.2), $\textrm{Cost}_{jt}$, as the instrument. Intuitively, cost is correlated with a brand’s price but not correlated with unobserved demand shocks, and it is a standard instrument for correcting price endogeneity in the demand estimation literature (Berry et al. 1995; Berry and Haile 2021).
To obtain control functions for two endogenous variables, we run the following two regressions and obtain the residuals; see (Aparicio et al. 2024) for a similar use case of this approach.
@@ 6.2 Price Endogeneity
[…]
@@
$+\alpha_{3}\mathcal{I}\{\textrm{BuyLastMonth}_{ijt}\}+\delta_{ij}+\lambda_{1}\textrm{CF\_Price}_{jt}+\lambda_{2}\textrm{CF\_Price\_cumPercOnlineSpend}_{ijt}+\tilde{\epsilon}_{ijt}$
@@
where $T_{i}$ is the set of purchase occasions associated with customer $i$. The parameter vector of interest is $\boldsymbol{\theta}=\{\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\lambda_{1},\lambda_{2},\{\delta_{ij}\}_{i=1,\ldots,N;j=1,\ldots,J}\}$. Note that this is a high-dimensional parameter vector – in addition to population parameters such as $\alpha$s and $\lambda$s, we also have customer-brand fixed-effects ($\delta_{ij}$s), which scale with the sample size.
@@ 6.3 Estimation
Our main estimation strategy follows the standard two-step approach used in control function approaches; see (Petrin and Train 2010) for details. However, the key additional challenge in our setting comes from the presence of the high-dimensional customer-brand fixed-effects. We now discuss both the overview of our estimation strategy and also provide a summary of the solution concept that we adopt to address the high-dimensional fixed-effects.
[…]
@@
Based on the Theorem 1 and proof from Chen et al. (2021), the iterative sequence $\boldsymbol{\theta}^{(k+1)}=\arg\max_{\theta}S(\boldsymbol{\theta};\boldsymbol{\theta}^{(k)})$, $k=0,1,\ldots,$ converges to a local, if not global, maximum of $\mathcal{L}(\boldsymbol{\theta})$.  [^8] For more details, please refer to the proof of Theorem 1 in Chen et al. (2021), which shows the convergence of this iterative estimator to the true parameters.
Equation (11) is a second-order Taylor expansion of the log-likelihood function at $\boldsymbol{\theta}^{(k)}$. The model parameters we are searching over in the iterative MM algorithm $(\boldsymbol{\theta}$, rather than $\boldsymbol{\theta}^{(k)})$, appear only in the third term. Because the only term in $S(\boldsymbol{\theta};\boldsymbol{\theta}^{(k)})$ that depends on $\boldsymbol{\theta}$ is the sum of squared residuals, maximizing $S(\boldsymbol{\theta};\boldsymbol{\theta}^{(k)})$ is equivalent to minimizing this least squares objective, and the coefficients can be obtained by regressing $\psi_{ijt}^{(k)}+h_{j}(\boldsymbol{\psi}_{it}^{(k)};\boldsymbol{y}_{it})$ on the vector of explanatory variables $\boldsymbol{x}_{ijt}=\{\textrm{Price}_{jt},\textrm{cumPercOnlineSpend}_{it},\textrm{Price}_{jt}\times\textrm{cumPercOnlineSpend}_{it},\mathcal{I}\{\textrm{BuyLastMonth}_{ijt}\},\textrm{CF\_Price}_{jt},\textrm{CF\_Price\_cumPercOnlineSpend}_{ijt},\delta_{ij}\}$ in the utility function. This linearization allows us to exploit standard panel data methods despite a large number of fixed effects $\delta_{ij}$.
@@
Since the estimated residuals from the first stage enter the estimation in the second stage, the standard errors in the second step estimation tend to be downward biased, although the parameter estimates themselves are consistent (Wooldridge 2010). Therefore, we use block bootstrapping with 500 replications, which treats individual customers as blocks. In each replication, we resample the same number of customers as in the original sample with replacement, re-estimate the demand model, and then obtain the bootstrapped standard errors for the demand parameters and price elasticities (Cameron 2005).

### v2

@@ 3.1 Setting
We use data from a leading pet supplies retailer in Brazil. This firm operates in a multichannel environment, with over 200 brick-and-mortar stores nationwide and a digital presence through both a mobile app and an e-commerce website. The retailer offers a broad selection of pet-related products, including food, medicines, hygiene and grooming items, toys, and accessories. To track consumer purchases across online and offline channels, the retailer uses Brazil’s National Identification Number, which provides a unique identifier for each customer. Online purchases - whether through the app or website - require customers to log into their accounts, making all such transactions traceable to individual IDs. For brick-and-mortar purchases, customers provide their ID at checkout, although it is not mandatory. According to the retailer, around a quarter of offline transactions are not associated with a customer ID, and this rate has remained stable throughout the study period. [^2] The fact that this rate has remained stable despite the fact that there was a spike in online shopping adoption during COVID-19 reassures us that the adoption of online shopping does not affect it.
@@ 3.2 Dataset Construction
[…]
Selection of Customers: Following the standard approach in the brand choice literature (Kim et al., 2002; Chintagunta et al., 2005; Dubé et al., 2010), for each category, we restrict our sample to only those consumers who only purchase the specific pack size of interest for that category (as defined earlier), without switching to any other pack sizes within the same category. This is so that we are able to fully capture these customers’ substitution patterns within the category. Next, given that our goal is to investigate the effect of consumers’ adoption of online shopping at a given retailer333“Adoption of online shopping” refers to a customer making his/her first online purchase with the focal retailer in any category, and not just one of the four focal categories. on their offline price sensitivity at that retailer’s physical stores, we need to ensure that we have sufficiently long pre- and post-adoption periods for each adopter in our sample to effectively measure price elasticities pre- and post-adoption. Therefore, we focus on customers who adopt online shopping with the retailer in the middle of our observation period, i.e., July 1st, 2019 – June 30th, 2020. So our 30-month data period is divided into three phases: a pre-adoption period (January 1st, 2019 – June 30th, 2019) when all adopters are in their pre-adoption phase, an adoption period (July 1st, 2019 – June 30th, 2020) during which time the adopters in our sample make their first online purchase, and a post-adoption period (July 1st, 2020 - June 30th, 2021), when all the adopters in our sample are in their post-adoption phase. To clarify, a given adopter’s pre- and post-adoption phases are defined as the months before and after they make their first online purchase. Non-adopters only shop offline through the entire data period. [^4] Note that we also verified with the retailer that the identified adopters and non-adopters in our data did not adopt online shopping before our data period begins.
@@
Estimation Sample Construction: For each customer in a given category (both adopters and non-adopters), we organize the data into a panel that captures purchase occasions and brand choices for individual customers over time. Specifically, we define purchase occasions at the customer-month level. If a customer visits a physical store of the retailer during a month and purchases in any category (not limited to the categories we focus on in this study), that customer-month is included as a purchase occasion for all four categories. If a customer makes no purchases in any category during the month, that customer-month is not included as a purchase occasion. In other words, a purchase occasion represents any customer-month during which we observe any offline interaction with the retailer (similar to week-store visit in Gordon et al., (2013)). [^5] Based on our definition of purchase occasion, if a customer visits multiple times in a month, it is treated as one purchase occasion. In our data, customers, by and large, do not purchase from more than one store in a month. More than 96% of customer-month-level purchase occasions involve purchases from a single store. When we process the choice data for each category, for the minor fraction of purchase occasions that involve more than one store, we label these customer-month purchase occasions with the store where the customer spent the most within that category during the month so that each customer-month level purchase occasion is associated with a single store. This approach of using purchase occasions to construct the estimation sample is standard in the demand estimation literature studying brand choice using household level data, since it allows researchers to focus on modeling brand choice and purchase incidence in a category conditional on a store visit, instead of modeling the decision of when to go to the store or which store to visit (Chib et al., 2004; Chintagunta et al., 2005; Dubé et al., 2010; Gordon et al., 2013). Following the standard practice in the literature, we do not model the channel choice decision explicitly for a couple of reasons – (a) channel choice can depend on multiple factors that we do not observe (e.g. transportation costs, distance to stores, consumer expectations of prices etc.); (b) we would need to impose a structure on the order of store choice, channel choice and brand choice and it is ex-ante not clear in which order consumers make these decisions. Given that our focus is on documenting the empirical effect of the adoption of online shopping on price sensitivity in offline brand choice decisions, modeling the arrival of purchase occasions is neither central nor necessary to answer our research question. Further, in $\S$8.2.4, we show that our results are robust to alternative definitions of purchase occasions.
For each purchase occasion, we identify the brand chosen by the consumer as the brand that has positive spend. [^6] We rarely observe customer-month purchase occasions that involve a purchase of more than one brand. All product categories have fewer than 0.86% of total purchase occasions (e.g., including those with multiple visits to a single store in a month, visits to multiple stores and a single visit) with multiple brand purchases. For these few purchase occasions with multiple brand purchases, we assign the brand with the largest spend for that customer during the month, similar to previous brand choice literature, e.g. Erdem and Keane, (1996). Some researchers have handled this issue differently, for instance, by having each brand chosen count as an independent single-brand purchase occasion (Krishnamurthi and Raj, 1988; Chib et al., 2004). Since the fraction of purchases is so small, we do not think this decision makes a material impact on our results. If the customer does not have positive spend on any brand during a purchase occasion, we label it as choosing the outside option (i.e., “no-purchase”).
@@ 3.3 Summary Statistics
[…]
@@ 4 Descriptive Analysis
We now present some descriptive analysis that examines if/how the adoption of online shopping with the retailer changes consumer behavior during their offline shopping trips at the retailer’s physical stores in the post-adoption period. Specifically, we focus on two variables that can capture changes in within-customer offline price sensitivity:
@@ 4.1 Price Paid
$\mathcal{I}\{\textrm{Adopter}_{i}\}\times\mathcal{I}\{t\geq T_{i}^{\textrm{adopt}}\}$ -0.0102 -0.0106∗∗ -0.0046 -0.0160
@@
We examine how the prices paid by adopters change after online adoption (compared to non-adopters). To that end, we focus on all offline purchase orders (made by adopters and non-adopters) that contain a focal category purchase and estimate the following two-way fixed-effects regression for each category at the customer-order level:
$\textrm{ln(Price\_Paid)}_{it}=\delta\times\mathcal{I}\{\textrm{Adopter}_{i}\}\times\mathcal{I}\{t\geq T_{i}^{\textrm{adopt}}\}+\textrm{Customer}_{i}+\textrm{YearMonth}_{t}+\epsilon_{it},$ (1)
@@ 4.2 Category Purchase Frequency
$\mathcal{I}\{\textrm{Adopter}_{i}\}\times\mathcal{I}\{t\geq T_{i}^{\textrm{adopt}}\}$ -0.0751∗∗∗ -0.0497∗∗∗ -0.0617∗∗∗ -0.0654∗∗∗
@@ 5 Firm’s Problem and Overview of Empirical Strategy
The primary objective of the firm is to quantify the causal effect of consumers’ adoption of its online channels on their offline price sensitivity at its physical stores, compared to consumers who shop exclusively offline. However, quantifying this impact presents both methodological challenges and limited managerial relevance for several reasons.
First, adopting the online channel is a consumer-level decision, and firms cannot randomly assign consumers to adopt the online channel or remain offline-only through field experiments. This makes it impossible to directly measure the Average Treatment Effect (ATE) of online adoption on consumers’ offline price sensitivity. Second, although firms could implement an intent-to-treat (ITT) design - such as randomly offering incentives to some offline-only consumers to encourage their online adoption - this approach still does not give us ATE given that some consumers would never adopt online  [^8] We do not obtain ATE under the ITT design, since the causal quantity obtained from 2SLS estimators represents the Average Causal Response, which is a weighted sum of individual treatment effects where weights are compliance scores; For more details, see Angrist and Imbens, (1995); Syrgkanis et al., (2019); Chen et al., (2024); Mummalaneni et al., (2025)..Third, even if a firm could somehow identify a significant effect of adopting online shopping on consumers’ offline price sensitivity, it cannot force consumers to choose a specific channel. Therefore, obtaining the ATE is not only impractical but also lacks actionable implications. Instead, the main quantity of interest for the firm is the change in offline price sensitivity among consumers who adopt online shopping compared to those who continue to shop exclusively offline i.e. the Average Treatment Effect on the Treated (ATT). Measuring this would help firms adjust their offline pricing strategies in response to their customers adopting online shopping. Thus, our goal is to identify the ATT of online adoption on offline price sensitivity.
@@
    First, we measure category-level own-price elasticities at the customer-month level: In $\S$6, we use a logit brand choice model where each consumer’s price sensitivity coefficient varies with the extent of their online spending and include customer-brand fixed effects. To address price endogeneity, we use the control function approach (Petrin and Train, 2010), using wholesale cost as an instrument. To efficiently estimate the model given a large number of customer-brand fixed effects, we implement a Minorization-Maximization (MM) algorithm for the logit model (Chen et al., 2022).
@@
    Next, we use the estimated elasticities in the first step to measure the change in offline price elasticities due to online adoption. In $\S$7, we use a staggered difference-in-differences approach (Callaway and Sant’Anna, 2021) to estimate the ATT of online adoption on offline price elasticities after accounting for differences in the timing of adoption between different adopter cohorts.
@@ 6.1 Choice Model and Utility Specification
[…]
@@
$+\alpha_{3}\mathcal{I}\{\textrm{BuyPrevOcca}_{ijt}\}+\delta_{ij}+\epsilon_{ijt}~~\forall j\in\{1,\ldots,J\}$
In this utility specification, $\textrm{Price}_{jt}$ is the price of brand $j$ during purchase occasion $t$. We allow a consumer’s price sensitivity coefficient to depend on their cumulative percentage of online spending, denoted as $\textrm{cumPercOnlineSpend}_{it}$. This variable represents the proportion of the customer’s total spending with the retailer that has occurred online up to the purchase occasion $t$ since their first transaction with the retailer. We use this continuous variable to allow for variation in the online adoption effect based on the extent of a consumer’s online spend. We also consider an alternate specification where we use a binary online adoption indicator in §8.2.1. The coefficient $\alpha_{0}$ represents the baseline price sensitivity for purely offline customers. The interaction coefficient $\alpha_{2}$ reflects how price sensitivity changes as consumers spend more online. A negative $\alpha_{2}$ would suggest that higher online spending share leads to greater offline price sensitivity. [^11] Interaction terms in logit and other nonlinear models are best interpreted via marginal effects or elasticities (Ai and Norton, 2003). The variable $\mathcal{I}\{\textrm{BuyPrevOcca}_{ijt}\}$ is an indicator that equals 1 if customer $i$ purchased brand $j$ at the previous purchase occasion t1t-1, and 0 otherwise, capturing state dependence via $\alpha_{3}$. The customer-brand fixed effects $\delta_{ij}$ capture consumer $i$’s time-invariant preference for brand $j$. These fixed effects capture systematic time-invariant consumer-level heterogeneity in brand preferences. However, including these high-dimensional fixed effects introduces some estimation challenges which we discuss in §6.3. The term $\epsilon_{ijt}$ is an idiosyncratic error term. Finally, the utility of choosing the outside option is normalized as $U_{i0t}=0+\epsilon_{i0t}$ where $\epsilon_{i0t}$ also follows an i.i.d. Type I extreme value distribution.
6.2 Price Endogeneity
A common challenge in estimating price elasticities with observational choice data is price endogeneity. This issue arises when the error term, $\epsilon_{ijt}$, contains some time-varying unobserved demand shocks that may be correlated with prices; e.g., if promotions or advertising are correlated with prices. To address this, we apply the control function approach for two endogenous variables, $\textrm{Price}_{jt}$ and $\textrm{Price}_{jt}\times\textrm{cumPercOnlineSpend}_{it}$ (Petrin and Train, 2010; Wooldridge, 2015). We use the weighted wholesale cost indices for each brand $j$ (constructed as described in $\S$3.2), $\textrm{Cost}_{jt}$, as the instrument. Intuitively, cost is correlated with a brand’s price but not correlated with unobserved demand shocks, and it is a standard instrument for correcting price endogeneity in the demand estimation literature (Berry et al., 1995; Berry and Haile, 2021).
@@ 6.2 Price Endogeneity
Here, $\textrm{Control}_{ijt}$ is a vector of control variables (e.g., $\textrm{cumPercOnlineSpend}_{it},\mathcal{I}\{\textrm{BuyPrevOcca}_{ijt}\},\textrm{Customer}_{i},$ $\textrm{Brand}_{j},\textrm{YearMonth}_{t}$). Next, following Petrin and Train, (2010), the original error term in the utility specification, $\epsilon_{ijt}$, can be decomposed as follows:
@@
$+\alpha_{3}\mathcal{I}\{\textrm{BuyPrevOcca}_{ijt}\}+\delta_{ij}+\lambda_{1}\textrm{CF\_Price}_{jt}+\lambda_{2}\textrm{CF\_Price\_cumPercOnlineSpend}_{ijt}+\tilde{\epsilon}_{ijt}$
@@
[…]
@@ 6.3 Estimation
Our main estimation strategy follows the standard two-step approach used in control function approaches; see (Petrin and Train, 2010) for details. However, the key additional challenge in our setting arises from the presence of high-dimensional customer-brand fixed effects. While high-dimensional fixed effects generally allow for more flexible substitution patterns (Jiang et al., 2025), doing so complicates the estimation procedure. We now discuss both the overview of our estimation strategy and also provide a summary of the solution concept that we adopt to address the high-dimensional fixed-effects.
First, we run OLS regressions for $\textrm{Price}_{jt}$ in Equation (4) and $\textrm{Price}_{jt}\times\textrm{cumPercOnlineSpend}_{it}$ in Equation (5), respectively. We then obtain the residuals, $\textrm{CF\_Price}_{jt}$ and $\textrm{CF\_Price\_cumPercOnlineSpend}_{ijt}$, which serve as control functions. Second, we include these residuals as additional regressors in the logit model (i.e., Equation (7)). In a standard setting, estimation can proceed as usual at this stage. However, since we have a large number of customer-brand fixed effects $(\delta_{ij})$, standard maximum likelihood estimation becomes computationally infeasible. To overcome this, we adopt a Minorization-Maximization (MM) algorithm for the logit model with fixed effects, an efficient estimation method proposed by Chen et al., (2022). The MM algorithm transforms the non-linear optimization problem (i.e., MLE estimation) into a sequence of simpler problems by iteratively constructing and maximizing a surrogate transfer minorization function $S(\cdot)$ that minorizes the log-likelihood (James, 2017). The transfer minorization function $S(\cdot)$ is less than the log-likelihood function everywhere except at the current best guess of the parameters, where it is equal. Chen et al., (2022) derives a transfer minorization function for the standard multinomial logit model in which there are no alternative-specific time-varying covariates in the utility function. In the theorem below, we extend their proof to cover a fully flexible logit model with covariates that may vary by individual, by alternative (brand), and over time. [^13] While our paper was the first to present this extension and proof for this specific case, in a recently posted update, Chen et al., (2025) present a similar extension. We then derive the corresponding transfer minorization function $S(\cdot)$ for the log-likelihood $\mathcal{L}$.
@@
Based on the Theorem 1 and proof from Chen et al., (2022), the iterative sequence $\boldsymbol{\theta}^{(k+1)}=\arg\max_{\theta}S(\boldsymbol{\theta};\boldsymbol{\theta}^{(k)})$, $k=0,1,\ldots,$ converges to a local, if not global, maximum of $\mathcal{L}(\boldsymbol{\theta})$. [^14] For more details, please refer to the proof of Theorem 1 in Chen et al., (2022), which shows the convergence of this iterative estimator to the true parameters.
Equation (11) is a second-order Taylor expansion of the log-likelihood function at $\boldsymbol{\theta}^{(k)}$. The model parameters we are searching over in the iterative MM algorithm $(\boldsymbol{\theta}$, rather than $\boldsymbol{\theta}^{(k)})$, appear only in the third term. Because the only term in $S(\boldsymbol{\theta};\boldsymbol{\theta}^{(k)})$ that depends on $\boldsymbol{\theta}$ is the sum of squared residuals, maximizing $S(\boldsymbol{\theta};\boldsymbol{\theta}^{(k)})$ is equivalent to minimizing this least squares objective, and the coefficients can be obtained by regressing $\psi_{ijt}^{(k)}+h_{j}(\boldsymbol{\psi}_{it}^{(k)};\boldsymbol{y}_{it})$ on the vector of explanatory variables $\boldsymbol{x}_{ijt}=\{\textrm{Price}_{jt},\textrm{cumPercOnlineSpend}_{it},\textrm{Price}_{jt}\times\textrm{cumPercOnlineSpend}_{it},\mathcal{I}\{\textrm{BuyPrevOcca}_{ijt}\},\textrm{CF\_Price}_{jt},\textrm{CF\_Price\_cumPercOnlineSpend}_{ijt},\delta_{ij}\}$ in the utility function. This linearization enables us to apply standard panel data methods despite the large number of fixed effects $\delta_{ij}$. Moreover, another advantage of using the MM algorithm to estimate the logit model is that for customer-brand pairs with few or no purchases, maximum likelihood estimation formally drives $\delta_{ij}$ toward -$\infty$, so that the predicted choice probability is exactly zero. In practice, this boundary solution is numerically unstable. The MM algorithm we employ circumvents this by iteratively solving a series of linearized regressions: never-chosen customer-brand pairs converge to large negative but finite estimates, which approximate the boundary solution while preserving numerical stability. Thus, the algorithm can include all $(i,j)$ pairs without manual dropping, effectively regularizing the fixed effects associated with rate or absent choices.
@@
Since the estimated residuals from the first stage enter the estimation in the second stage, the standard errors in the second step estimation tend to be downward biased, although the parameter estimates themselves are consistent (Wooldridge, 2010). Therefore, we use block bootstrapping with 500 replications, which treats individual customers as blocks. In each replication, we resample the same number of customers as in the original sample with replacement, re-estimate the demand model, and then obtain the bootstrapped standard errors for the demand parameters and price elasticities (Cameron, 2005).